# Enhanced thermoelectric effects in a driven one-dimensional system

### v1

@@ V Conclusions
In the present work we investigated the effects of an AC driving on the thermoelectric properties in the framework of Floquet theory. Two simple models are studied in one-dimensional space, both including a delta-function like potential whose strength oscillates in time (the AC driving). In the first model, the system has left-to-right symmetry, which can be seen as a homogeneous highly-doped 1D semiconductor, while in the second model this symmetry is broken by introducing a step function potential, which may describe a inhomogeneous semiconducting interface. We assume that the electron density in the system is very low such that the AC driving frequency can be as large as fractions of the Fermi energy and/or comparable with the semiconducting gap.
Our results for the first model showed that the driving could strongly enhance the Seebeck coefficient, especially at low temperatures, by increasing the strength of the AC component of the delta-like potential with respect to the static one. In particular, we found that the relative change of the Seebeck coefficient with respect to the static condition could reach 200% at large frequencies, where the corresponding energy significantly exceeds the thermal energy, i.e. in the quantum regime. In the second model, we focused on the thermoelectric Onsager coefficient, finding that its relative change with respect to the static case could be enhanced by a large amount at low temperatures, especially when the chemical potential is within the gap of the semiconductor, where the current is negligible in static conditions. In the second model, where left-to-right symmetry is broken, we note that additional current contributions arise from a non-adiabatic pumping effect. This effect has the potential to partially obscure the enhancement of thermoelectricity observed in the symmetric case. Despite this, by isolating the terms linear in the temperature difference, it is still possible to single out the impact of photon-assisted processes on thermoelectricity. The second model offers a potentially useful analogy for phenomenology at a normal metal/superconductor interface, where the superconducting gap arises from pairing.
To conclude, our results demonstrate that external driving (i.e. coherent photon-assisted processes) could be used to tune and enhance the thermoelectric capabilities of low-electron-density nanodevices.
@@ VI Acknowledgements
We acknowledge the Royal Society through the International Exchanges between the UK and Italy (Grants No. IEC/R2/212041). F.T. and A.B. acknowledge the MUR-PRIN 2022—Grant No. 2022B9P8LN-(PE3)-Project NEThEQS ”Non-equilibrium coherent thermal effects in quantum systems” in PNRR Mission 4-Component 2-Investment 1.1 ”Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)” funded by the European Union-Next Generation EU and CNR project QTHERMONANO. A.B. acknowledges discussions with F. Giazotto and A. Jordan.
@@ Appendix A Methods to calculate the transmission coefficients within the Floquet theory
As mentioned in Sec. III of the main text, to obtain the vector of transmission coefficients $\vec{\xi}$ in the Floquet theory, one has to treat the infinite linear equations $\mathbf{M}\cdot\vec{\xi}=\vec{B}$, representing the set of Eqs. (13), with $\mathbf{M}$ an infinite-sized known matrix and $\vec{B}$ a known vector related to the amplitude of the incident wave. The matrix $\mathbf{M}$ can be decomposed into $\mathbf{M}=\mathbf{\Lambda}+u_{1}\mathbf{M}_{1}$, in which the matrix $\mathbf{\Lambda}$ is a diagonal matrix [with elements $(...,\Lambda_{1},\Lambda_{0},\Lambda_{-1},...)$], and the matrix $\mathbf{M}_{1}$ is zero on the diagonal. We can use two approximate methods to treat the problem. The first one is perturbative: $\vec{\xi}$ can be calculated order by order by assuming that $u_{1}$ is a small parameter. The second method is to truncate the matrix $\mathbf{M}$ into a finite dimension and then solve the truncated matrix equation (finite dimensional). In practice, the latter approach has to be done numerically. Both methods are accurate to a certain order (power) in $v=u_{1}/u_{0}$, and their numerical results clearly merge when the order increases.
@@
Figure 8: Transmission function $\mathcal{T}(E)$: comparison between the perturbation result (orange curve) and the truncation result (blue curve), with the parameters $\eta=0.1$ and $v=0.4$. As for the truncation, we keep 5 positive modes and 5 negative modes, i.e. $n=-5,-4,...,5$, in total 11 components. The black dashed curve is the static case. The vertical dotted line indicates the energy of the AC driving with $\hbar\omega=0.2\mu$.
Interestingly, the occurrence of the kink structure in the transmission function around $E=\hbar\omega$ (especially the sharp drop on the right hand side of kink, i.e. for $E>\hbar\omega$) can be explained using the perturbative method as follows. If the amplitude of the AC signal is small, i.e. $v=u_{1}/u_{0}\ll 1$, we can express the transmission function as $\mathcal{T}(E)=|\xi^{(0)}_{0}+u_{1}^{2}\xi^{(2)}_{0}|^{2}+u_{1}^{2}|\xi^{(1)}_{1}|^{2}+u_{1}^{2}|\xi^{(1)}_{-1}|^{2}$, in which the terms with higher powers than $u_{1}^{2}$ are neglected. Inserting the results obtained above, we have $\mathcal{T}(E)=(k_{0}^{2}/|\Lambda_{0}|^{2})\mathcal{W}(E)$ with $\mathcal{W}=\mathcal{W}_{0}+\mathcal{W}_{1}+\mathcal{W}_{-1}$ in which
@@
with $\mathcal{N}=(1-\varphi_{0}^{2})/(\varphi_{0}(1+\varphi_{0}^{2}))$. In Eq. (23), terms with higher power than $v^{2}$ are omitted, and $\varphi_{0}$ is the limit value of $\phi_{0}$ when $\epsilon\to 0$. By comparing with the notations introduced in Sec. III, one finds that $\varphi_{0}^{2}=(\hbar\omega/E_{F})/\eta^{2}$. For the case with $\eta=0.1$ and $\hbar\omega=0.2\mu$, we have $\mathcal{N}<0$, and therefore $\mathcal{T}(E)$ decreases with $E$, when $E$ is a little bigger than $\hbar\omega$. Furthermore, $\delta$ has a strong dependence on $E$, when $E$ is near to $\hbar\omega$ from the right side, which explains the sharp peak structure at $E=\hbar\omega$.
@@ Appendix B Pumped current
[…]
@@ C.1 Single barrier
[…]
@@
Consequently, when the temperature is sufficiently high for $f^{\prime}(E)$ to encompass this sharp feature, the integrand reduces in value. This leads to a decrease in $S/T$ starting from $k_{B}T\simeq 0.02\mu$ for the green curve. In the case $\hbar\omega=0.8\mu$ (purple curve), the effect of the kink structure on $S$ is much weaker, but still causing a decrease of $S/T$ starting from $k_{B}T\simeq 0.04\mu$. In the other cases involving the red ($\hbar\omega=0.5\mu$) and blue ($\hbar\omega=0.2\mu$) curves, the kink structure at $E=\hbar\omega$ gradually moves away from the chemical potential. This leads to a behavior of $\mathcal{L}(E)$ that closely resembles the static case (see Fig. 10). As a result, the Seebeck coefficient behaves similarly to the static scenario, following the black dashed curve and increasing at higher temperatures.
@@ C.2 Single barrier with a step
In order to understand the temperature dependence of $L_{\rm eh}/T^{3}$, we follow the approach of the previous subsection and introduce the auxiliary function $\mathcal{L}(E)$, defined by Eq. (25), in which the transmission function $\mathcal{T}(E)$ is replaced by the averaged transmission $\bar{\mathcal{T}}(E)$. Fig. 11 shows the plot of function $\mathcal{L}(E)$ with non-zero potential step. The value of $\mathcal{L}(\mu)$ is very close to the value of the derivative of $\bar{\mathcal{T}}(E)$ at the chemical potential, i.e. $\bar{\mathcal{T}}^{\prime}(\mu)$. A key feature of the curves of $\mathcal{L}(E)$ in Fig. 11 are the peaks at $E=\Delta$. As mentioned in the previous subsection, at $T=0$, the function $f^{\prime}(E)$ is a delta function centered at $E=\mu$. When temperature increases, the function $f^{\prime}(E)$ broadens and the integration in Eq. (8) averages the function $(E-\mu)^{2}\mathcal{L}(E)$ over an interval around $E=\mu$. If the broadening includes the peak structure of $\mathcal{L}(E)$, the value of $L_{\rm eh}/T^{3}$ increases. However, at higher temperatures, when the averaging interval extends beyond the peak, the contribution from the peak diminishes, leading to a decrease in $L_{\rm eh}/T^{3}$.

### v2

@@ V Conclusions
In the present work, we investigated the effects of an AC driving on the thermoelectric properties in the framework of Floquet theory. Two simple models are studied in one-dimensional space, both including a delta-function like potential whose strength oscillates in time (the AC driving). In the first model, the system exhibits left-to-right symmetry, which can be viewed as a homogeneous, highly doped 1D semiconductor. In contrast, the second model breaks this symmetry by introducing a step function potential, which may describe an inhomogeneous semiconducting interface. We assume that the electron density in the system is very low, such that the AC driving frequency can be as large as fractions of the Fermi energy and/or comparable with the semiconducting gap.
Our results for the first model showed that driving could strongly enhance the Seebeck coefficient, especially at low temperatures, by increasing the strength of the AC component of the delta-like potential with respect to the static one. In particular, we found that the relative change of the Seebeck coefficient with respect to the static condition could reach 200% at large frequencies, where the corresponding energy significantly exceeds the thermal energy, i.e. in the quantum regime. In the second model, we focused on the thermoelectric Onsager coefficient, finding that its relative change with respect to the static case could be significantly enhanced at low temperatures, especially when the chemical potential is within the semiconductor’s gap, where the current is negligible in static conditions.
In the second model, where left-to-right symmetry is broken, we note that additional current contributions arise from a non-adiabatic pumping effect. This effect has the potential to partially obscure the enhancement of thermoelectricity observed in the symmetric case. Despite this, by isolating the terms linear in the temperature difference, it is still possible to single out the impact of photon-assisted processes on thermoelectricity. The second model not only describes inhomogeneous semiconductors but also offers a potentially useful analogy for phenomenology at a normal metal/superconductor interface, where the superconducting gap arises from the pairing of electrons. Such an analogy appears relevant for the field of superconducting quantum technologies, where superconductors are operated at cryogenic temperatures in the high-frequency (quantum) domain.
To conclude, our results demonstrate that external driving (i.e. coherent photon-assisted processes) can be used to tune and enhance the thermoelectric capabilities of low-electron-density nanodevices. For the future, it would be very interesting to investigate the effect of high-frequency driving on a resonant system, where thermoelectric properties are already excellent in the static situation.
@@ VI Acknowledgements
We acknowledge the Royal Society through the International Exchanges between the UK and Italy (Grants No. IEC/R2/212041). F.T. and A.B. acknowledge the MUR-PRIN 2022—Grant No. 2022B9P8LN-(PE3)-Project NEThEQS “Non-equilibrium coherent thermal effects in quantum systems” in PNRR Mission 4-Component 2-Investment 1.1 “Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)” funded by the European Union-Next Generation EU and CNR project QTHERMONANO. A.B. acknowledges discussions with F. Giazotto and A. Jordan.
@@ Appendix A Methods to calculate the transmission coefficients within the Floquet theory
[…]
@@
Figure 8: Transmission function $\mathcal{T}(E)$: comparison between the perturbation result (orange curve) and the truncation result (blue curve), with the parameters $\eta=0.1$ and $v=0.4$. As for the truncation, we keep 5 positive modes and 5 negative modes, i.e. n=5,4,,5n=-5,-4,...,5, with a total of 11 components. The black dashed curve is the static case. The vertical dotted line indicates the energy of the AC driving with $\hbar\omega=0.2\mu$.
Interestingly, the occurrence of the kink structure in the transmission function around $E=\hbar\omega$ (especially the sharp drop on the right-hand side of the kink, i.e. for $E>\hbar\omega$) can be explained using the perturbative method as follows. If the amplitude of the AC signal is small, i.e. $v=u_{1}/u_{0}\ll 1$, we can express the transmission function as $\mathcal{T}(E)=|\xi^{(0)}_{0}+u_{1}^{2}\xi^{(2)}_{0}|^{2}+u_{1}^{2}|\xi^{(1)}_{1}|^{2}+u_{1}^{2}|\xi^{(1)}_{-1}|^{2}$, in which the terms with higher powers than $u_{1}^{2}$ are neglected. Inserting the results obtained above, we have $\mathcal{T}(E)=(k_{0}^{2}/|\Lambda_{0}|^{2})\mathcal{W}(E)$ with $\mathcal{W}=\mathcal{W}_{0}+\mathcal{W}_{1}+\mathcal{W}_{-1}$ in which
@@
with $\mathcal{N}=(1-\varphi_{0}^{2})/(\varphi_{0}(1+\varphi_{0}^{2}))$. In Eq. (23), terms with higher power than $v^{2}$ are omitted, and $\varphi_{0}$ is the limit value of $\phi_{0}$ when $\epsilon\to 0$. By comparing with the notations introduced in Sec. III, one finds that $\varphi_{0}^{2}=(\hbar\omega/E_{F})/\eta^{2}$. For the case with $\eta=0.1$ and $\hbar\omega=0.2\mu$, we have $\mathcal{N}<0$, and therefore $\mathcal{T}(E)$ decreases with $E$, when $E$ is a little bigger than $\hbar\omega$. Furthermore, $\delta$ has a strong dependence on $E$, when $E$ is near $\hbar\omega$ from the right side, which explains the sharp peak structure at $E=\hbar\omega$.
@@ Appendix B Pumped current
$J_{e}=\frac{e}{h}\int_{-\infty}^{\infty}dE\left[\mathcal{T}_{RL}(E)f_{L}(E)\right.-\left.\mathcal{T}_{LR}(E)f_{R}(E)\right],$ (24)
@@ C.1 Single barrier
In order to understand the behavior of the Seebeck coefficient in Fig. 3, we need to focus on the numerator of Eq. (6), since the denominator, which represents the conductance, remains virtually temperature-independent in the range of temperatures considered (varying by only about 1%). The numerator of the Seebeck coefficient, proportional to $L_{\rm eh}$, can be expressed as $\int dE(E-\mu)^{2}\mathcal{L}(E)f^{\prime}(E)$, where we introduce the auxiliary function
@@
Consequently, when the temperature is sufficiently high for $f^{\prime}(E)$ to encompass this sharp feature, the integrand reduces in value. This leads to a decrease in $S/T$ starting from $k_{B}T\simeq 0.02\mu$ for the green curve. In the case $\hbar\omega=0.8\mu$ (purple curve), the effect of the kink structure on $S$ is much weaker, but still causes a decrease of $S/T$ starting from $k_{B}T\simeq 0.04\mu$. In the other cases involving the red ($\hbar\omega=0.5\mu$) and blue ($\hbar\omega=0.2\mu$) curves, the kink structure at $E=\hbar\omega$ gradually moves away from the chemical potential. This leads to a behavior of $\mathcal{L}(E)$ that closely resembles the static case (see Fig. 10). As a result, the Seebeck coefficient behaves similarly to the static scenario, following the black dashed curve and increasing at higher temperatures.
@@ C.2 Single barrier with a step
In order to understand the temperature dependence of $L_{\rm eh}/T^{3}$, we follow the approach of the previous subsection and introduce the auxiliary function $\mathcal{L}(E)$, defined by Eq. (25), in which the transmission function $\mathcal{T}(E)$ is replaced by the averaged transmission $\bar{\mathcal{T}}(E)$. Fig. 11 shows the plot of the function $\mathcal{L}(E)$ with a non-zero potential step. The value of $\mathcal{L}(\mu)$ is very close to the value of the derivative of $\bar{\mathcal{T}}(E)$ at the chemical potential, i.e. $\bar{\mathcal{T}}^{\prime}(\mu)$. A key feature of the curves of $\mathcal{L}(E)$ in Fig. 11 is the peaks at $E=\Delta$. As mentioned in the previous subsection, at $T=0$, the function $f^{\prime}(E)$ is a delta function centered at $E=\mu$. When temperature increases, the function $f^{\prime}(E)$ broadens and the integration in Eq. (8) averages the function $(E-\mu)^{2}\mathcal{L}(E)$ over an interval around $E=\mu$. If the broadening includes the peak structure of $\mathcal{L}(E)$, the value of $L_{\rm eh}/T^{3}$ increases. However, at higher temperatures, when the averaging interval extends beyond the peak, the contribution from the peak diminishes, leading to a decrease in $L_{\rm eh}/T^{3}$.